The length of a rectangle is four times its width.
If the area of the rectangle is 324m^2 , find its perimeter

Answer by mananth(16946)

width -------x
length -----------4x
area = 4x^2 =324
x^2=81
x= 9m the width
length = 9*4 = 36
perimeter = 36+36+9+9=90m